OpenStudy (anonymous):

I need help with my homework please. It is solving problems using rational equations in word problem form

OpenStudy (king):

So, tell the problems.....

OpenStudy (king):

Post them as a question

OpenStudy (kira_yamato):

We'll help you as far as we can...

OpenStudy (anonymous):

alana and her father begin skiing the same cross-country trail at ther same time. alana averages 9mph and finishes in 0.25hrs sooner. her father averages 6mph. how long is the trail?

OpenStudy (anonymous):

I have to make a key, an equation, and show my work but I can't even understand to even start

OpenStudy (kira_yamato):

alana and her father begin skiing the same cross-country trail at ther same time. alana averages 9mph and finishes in 0.25hrs sooner. her father averages 6mph. how long is the trail? Let the length of trail be x miles. Time taken by Alana = x/9 Time taken by father = x/6 \[\frac{x}{6} - \frac{x}{9} = 0.25\]

OpenStudy (kira_yamato):

Just solve the equation I provided for x, and you will get it

OpenStudy (anonymous):

ok

OpenStudy (kira_yamato):

3x/54 = 0.25 x = 4.5 miles

OpenStudy (anonymous):

ok you lost me how did u get that?

OpenStudy (king):

take lcm and then 9x-6x/54=0.25 3x=0.25*54 x=0.25*54/3 x=4.5

OpenStudy (kira_yamato):

I've created the var. x to rep. the length of the path. So the time taken by Alana = x/9 and that by father = x/6 Because time = dist. / ave. speed It is given (and taken for granted) that time interval b/w Alana and Father is 0.25 hours. So x/6 - x/9 = 0.25 (QED)

OpenStudy (king):

Let the length of trail be x miles. Here time=distance/speed So, thats why Kira took time as the following Time taken by Alana = x/9 Time taken by father = x/6 Given that Alana finished 0.25 hrs earlier than her father. So,x/6-x/9=0.25 take lcm and then 9x-6x/54=0.25 \[3x=0.25\times54\] \[x=\left[ 0.25\times54 \right]\div3\] x=4.5